OpenStudy (anonymous):

Need assurance and a little help! A parabola is given parametrically by x=at^2 , y= 2at. If P is any point on the parabola, let F be the foot of the perpendicular from P onto the axis of symmetry. Let G be the point where the normal from P crosses the axis of symmetry. Prove that FG=2a

OpenStudy (anonymous):

x' = 2at y'= 2a ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then 2a/2at = 1/t

OpenStudy (anonymous):

Use the equation of Normal \[y - 2at = -\frac{1}{\frac1t}(x - at^2) \]

OpenStudy (anonymous):

Which is y=(1/t)x + at

OpenStudy (anonymous):

??No

OpenStudy (anonymous):

Normal is perpendicular to tangent. \[m_{t} m_n = -1\]

OpenStudy (anonymous):

OH, Ok.

OpenStudy (anonymous):

Axis of symmetry is the X-axis.

OpenStudy (anonymous):

How do you get that in the form y=mx+b?

OpenStudy (dumbcow):

isn't F and G the same since the normal is also perpendicular to P and both are meeting at x-axis ?

OpenStudy (anonymous):

I am confused now :/

OpenStudy (anonymous):

I'm not sure.. :/

OpenStudy (anonymous):

dumbcow's right I think

OpenStudy (dumbcow):

no guys what i really meant was i think theres a typo

OpenStudy (dumbcow):

maybe it should be F is where tangent line meet x-axis ?

OpenStudy (anonymous):

That's what it says in the book... (I mean what I wrote)

OpenStudy (anonymous):

Normal:\[y - 2at = -xt + at^3\]\[y + xt = at^3 + 2at\]\(y = 0\)\(\implies x = at^2 + 2at\) Tangent: \[y = \frac{1}{t}x +at\]\(y = 0\) \(\implies x = -at^2\)

OpenStudy (anonymous):

Why at^3 and not at^2?

OpenStudy (anonymous):

\[y - 2at = -t(x - at^2)\]

OpenStudy (anonymous):

The equation of Normal

OpenStudy (anonymous):

Ok, and how do you carry on from there?

OpenStudy (anonymous):

I don't know, the question stops making sense from here.

OpenStudy (anonymous):

hmmm Ok....

OpenStudy (dumbcow):

lol